Rationalize the two term denominator. 40. 1-V3 15 41. 1+73
Answer #1

Sokarton : is ISB In the Rationalize we remove the root in denominator = 5 x 1+ 15 = 5(1+13) (1)²-1(52 : 02 62 = (a+b) (9-b)tity E + J5 S_x 1-13 | 1+ 3 1-15 JS (1-5) (1-15) . als = (a+b)(a-b) > JS (15) 1-3 | SIS 15 Answer